OpenStudy (anonymous):

A softball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 5. h = height t = time in seconds Find the maximum height of the ball. a) 12.2 feet b) 13.9 feet c) 8.3 feet d) 7.6 feet

OpenStudy (anonymous):

find the vertex, the max height is the second coordinate of the vertex if this is for algebra find the 1st derivative , set it equal to 0 if this is calculus, then use 1st derivative test to justify your result

OpenStudy (anonymous):

i really dnt know how to do this

OpenStudy (anonymous):

I am questioning those answer choices

OpenStudy (anonymous):

plot the function in a graphing utility and you will see he answer is 21 feet, but I do not see that answer, are you sure you copied the function correctly?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[h(t)=-16t^2+32t+5\]

OpenStudy (anonymous):

is this for algebra or calculus class?

OpenStudy (anonymous):

for algebra class

OpenStudy (anonymous):

those answer choices are wrong, the answer is 21 feet

OpenStudy (anonymous):

you can complete the square to rewrite in "vertex" form to see the vertex of the parabola or use that little formula that comes from completing square, it would be for \[f(x)=Ax^2+Bx+C\] form the vertex is given by \[\left(\frac{B}{2A},f\left(\frac{B}{2A}\right)\right)\]

OpenStudy (vincent-lyon.fr):

I agree with @ebbflo: answer is 21 feet.